OpenStudy (anonymous):

name the coordinates of the vertex and the focus of the parabola... 3y^2-8x-18y-5=0.. do this by completeing the square. show work please!!!! thanks!!!!!

OpenStudy (anonymous):

where did you get 14 on the right side of the equation?

OpenStudy (campbell_st):

by completing the square... to get the perfect square I needed to add 9 to both sides on the equation,,,, but I just realised the mistake...

OpenStudy (anonymous):

ya for some reason I thought id have to multlpy 3and 9 and add it to 5 to get 32

OpenStudy (campbell_st):

ok... the vertex form I use is \[(y - k)^2 = 4a(x - h)\] vertex is (h, k) and a is the focal length rewrting you get \[3y^2 - 18y = 8x + 5\] add 27 to both sides of the equation will allow you to complete the square in y \[3y^2 - 18y + 27 = 8x + 5 + 27\] which becomes \[3(y^2 - 6y + 9) = 8x + 32\] factoring and you get \[3(y - 3)^2 = 8(x + 4)\] divide both sides by 3 \[(y - 3) = 4 \times \frac{2}{3}(x + 4)\] so the vertex is (-4, 3) and the focal length is 2/3 which means the focus is (-4 + 2/3, 3) hope it helps

OpenStudy (anonymous):

@campbell_st how would I graph this..

OpenStudy (campbell_st):

well it looks like